OpenStudy (anonymous):

how to differentiate -cos x^3 x^2??

OpenStudy (anonymous):

use product rule

OpenStudy (anonymous):

please write with brackets. Cant really tell what is what

OpenStudy (anonymous):

u(x)P(x)=u'(x)p(x)+u(x)p'(x) in you in your case you may set -cosx^3 as u(x) and p(x) as x^2

OpenStudy (anonymous):

so when differentiate -cos^3, we get -sin^3?? is it?

OpenStudy (anonymous):

is the equation supposed cos(x^3) or (cosx)^3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\cos ^{3}\]

OpenStudy (anonymous):

which one you are still not clear

OpenStudy (anonymous):

-cosx^3 or -cos(x^3) ?

OpenStudy (anonymous):

\[\cos ^{3}=-\sin ^{3}\] is it true?

OpenStudy (anonymous):

which one of these two the first one the second one?

OpenStudy (anonymous):

no, im answering the ques. i asked what is the answer after defferentiate \[\cos ^{3}\],, so is it true that after differentiate i get \[-\sin ^{3}\]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

because when -3cosx^2

OpenStudy (anonymous):

hey what happened to you

OpenStudy (anonymous):

i dont get wht ur saying

OpenStudy (anonymous):

-cosx^3 = sinx^3

OpenStudy (anonymous):

i have check and the answer actually -3 cos x^2 (-sin x),,, but tanx btw